how many gal. of a 65% saline solution must be mixed with 8gal. of pure water to make a 25% solution ?
Answer by stanbon(75887)

how many gal. of a 65% saline solution must be mixed with 8gal. of pure water to make a 25% solution ?
----
Equation:
salt + salt = salt
0.65x + 0*8 = 0.25(x+8)
---
Multiply thru by 100 to get:
65x + 0 = 25x + 25*8
---
40x = 25*8
x = (1/5)(25
x = 5 gallons (amt. of 65% solution needed)
==================
Cheers,
Stan H.